the power of a lens is +2.0 D. Find its focal length and state what kind of lens it is.
1
Expert's answer
2012-09-06T11:57:55-0400
D = +2.0 D, D = 1 / f.
Focal length f = 1 / D = 1 / 2.0 = 0.5 m
As D >
0, this is positive (or converging) lens.